Okay. I've read through the other threads that pertain to this part on the machine. Has anyone actually solved this problem on the Discs of Tron Environmental?
We are working on a EDOT that is a cut version. We pulled the monitor out yesterday and capped it. After we replaced the monitor and the boards, we noticed that one of the red lights was stuck in the "ON" position. The sequence still chases when Sark is hit but that one light on both sides stays on.
Any clues or recommendations?
We are working on a EDOT that is a cut version. We pulled the monitor out yesterday and capped it. After we replaced the monitor and the boards, we noticed that one of the red lights was stuck in the "ON" position. The sequence still chases when Sark is hit but that one light on both sides stays on.
Any clues or recommendations?
